What is the voltmeter reading across a 20-ohm resistor connected in series with a 10-ohm resistor and a 60-volt power source?


1
Expert's answer
2021-05-06T19:16:54-0400

We need to calculate the reading of an ideal voltmeter which is connected in parallel combination of resistance "20\Omega"

Concept Used : Reading of voltmeter is same as potential difference between its two terminals.

Assumption : Resistance of and ideal voltmeter is very large so that it will not draw any current from the main circuit and gives more accurate value of potential difference and current passing through it can be taken as zero.

Given : Voltage of external power source "= 60V"

Resistance "R_1""= 10\Omega" and "R_2 = 20\Omega"

Law used : Ohm's Law

From diagram,

Both the resistance are in series combination to equivalent resistance will be "30\Omega" .

Current passing the them is given by Ohm's law

"I=\frac{V}{R_1+R_2} = \frac{60V}{30 \Omega} = 2A"

from Ohm's law

Potential difference across "R_2 = IR_2 = 2A \times20\Omega" "= 40V"

So reading of voltmeter will be same as potential difference across "R_2".

Final Answer : "40V"